OpenStudy (anonymous):

Solve 5x + 2y = 7 10x + 4y = 14

OpenStudy (anonymous):

what is 7 10x?

OpenStudy (anonymous):

10 x begins a new line on the bottom it looks like this: Solve: 5x+2y=7 10x+4y=14

Parth (parthkohli):

If you multiply the first equation by 2, then you get equations like \(10x+4y=14 \qquad 10x+4y=14\) What do you think the result is when we have two same equations?

OpenStudy (anonymous):

(1, 1) (2, ) No Solution Infinitely Many Solutions

Parth (parthkohli):

Can you answer my question I asked you right there? :)

OpenStudy (anonymous):

They cancel out . . . . No solution ?

Parth (parthkohli):

No, actually if we have two same equations, we have \(\textbf{infinitely many solutions}\).

Parth (parthkohli):

Remember that :)

OpenStudy (anonymous):

Thank you so much .

OpenStudy (anonymous):

x + 3y = 6 2x – y = 10